# Skull-to-Face: Anatomy-Guided 3D Facial Reconstruction and Editing

### v1

@@ 4.1 Dataset and Error Metric
$E=\frac{1}{d|G|}\sum_{v\in G}\leavevmode\nobreak\ \underset{u\in F}{\min}\leavevmode\nobreak\ \|v-u\|_{2},$ (7)
@@
4.2 Initial Face Generation
@@
4.3 Evaluation of Tissue Depth Distribution
4.3.1 TDD-Global Model
@@
In the second experiment, we modified the face shapes of celebrities to show the ability of 3D face adjustment by the tissue depth model. Fig. 6 shows the results of three celebrities: Christian Bale, Adele Adkins, and Ling Jia. The initial 3D face of a celebrity was generated by the DECA model. We then used tissue depths of the pretrained TDD-Global model to estimate proxy skull landmarks (Sec. 3.2). We adjust $C$ to sample various valid combinations of tissue depths (e.g. from thin to fat) and then compute the target facial landmarks. Finally, the proposed anatomy-guided face adaptation is applied to adapt the initial 3D face to the new facial landmarks. The modified face is compared with the picture of the same person in a different face shape.
@@ 4.3.2 TDD-Regional Model
Figure 8: Validating the influence of initial faces on the final results. (a) are the query skulls and facial landmarks. (b) is the synthesized portrait image. (c) are the initial 3D faces with texture and the reconstruction errors. (d) are our facial reconstructions, which and fit the landmark constraints well. The final reconstruction errors are similar: 0.899% (top) and 0.925% (bottom). They are both close to the GT face (e).
@@
4.4 Evaluation of Skull-to-Face Reconstruction
4.4.1 Influence of Initial Faces on Final Results
We chose the average tissue depths to compute the extended landmarks as facial constraints. Fig. 8 shows that with face refinement, our face adaptation model can effectively modify the face shape to align with the facial landmarks. With facial landmark constraints, the two initial faces, after adaptation, converge to a similar face geometry. Their geometries still exhibit subtle differences inherited from the initial faces (see error maps in (d)).
@@ 4.4.1 Influence of Initial Faces on Final Results
We also conducted a user study to assess the convergence quality of the final facial reconstruction results from different initial faces. For each question, four initial faces are randomly synthesized. Three of them are used to reconstruct faces from Skull A, and the other one is used to reconstruct a face from Skull B. Users are required to identify the face geometries amongst the reconstructed faces and select the unique face that is reconstructed from Skull B. The detailed explanation and results of the user study are included in the Supplementary. The high average correctness of this user study indicates that our model can adapt different initial faces and converge to a face shape space that is related to the query skull, which is distinct from the reconstructed faces that are from the other skulls.
Table 1: Quantitative comparisons of facial reconstructions with existing methods. The distance between two ears of the groundtruth skin is normalized to 1.
@@
Table 2: Reconstruction errors before vs after adaptation and loss functions. Ablation Study Mean Recon. Mean Recon. (%) (millimeter) 1. Before adaptation 2.631 5.262   (from DECA) 2. w/ $\mathcal{L}_{proj}$ 2.273 4.547 3. w/ $\mathcal{L}_{lmk}$ 1.836 3.672 4. w/ $\mathcal{L}_{lmk}+\mathcal{L}_{proj}$ 1.758 3.576 5. After adaptation 1.732 3.464 ($\mathcal{L}_{lmk}+\mathcal{L}_{proj}+\mathcal{L}_{sym}$)
4.4.2 Comparisons with SOTA Facial Reconstruction Methods
@@
Tab. 4.4.1 shows the quantitative comparisons of facial reconstructions. Our method used synthesized portraits as initial faces. Ours (avg. tissue) used the average tissue depths from TDD-Global to compute the facial landmarks and then adapted the initial faces to the final results. Ours (best tissue) utilized three types of tissue depths thin, normal, and fat from TDD-Global as shown in Fig. 3. We selected the smallest reconstruction result as the final face. Our method outperforms existing methods in all metrics, largely due to the effective adaptation of our loss function to geometric constraints. The CT scan resolution is 0.25mm/pixel. The average of the reconstruction error is 3.464 millimeters.
@@
To evaluate the distribution bias of training/test set, we used 5-fold cross-validation to split the dataset into training (80%) and testing (20%). The average reconstruction errors on the five testing sets are: 1.761%, 1.454%, 1.828%, 1.575%, and 1.665%. The results are consistent with Table 4.4.1.
4.4.3 Validation on Real-world Facial Reconstruction
@@
4.5 Ablation Study on Face Adaptation
We conducted an ablation study to evaluate the impact of various loss functions via three settings: (1) using only $\mathcal{L}_{lmk}$, (2) using $\mathcal{L}_{lmk}$ + $\mathcal{L}_{proj}$, and (3) our complete loss. Tab. 2 reports the average reconstruction errors. We computed the errors in the physical metric (millimeter) by the average face width (20cm). The projection loss $\mathcal{L}_{proj}$ aids in reducing the distance between the reconstructed face and landmark constraints. The symmetric loss $\mathcal{L}_{sym}$ enhances the face realism by promoting symmetry. The second and third rows show the face adaptation results of the same skull, using two different initial faces.
5 Conclusions
We proposed a novel anatomy-guided 3D face reconstruction approach that can reconstruct faces from a given skull using both biological profiles of the skull and the statistical tissue depth information defined on physiological landmarks. Our pipeline includes three main designs, the biological profile-based 3D face initialization, the facial landmark prediction based on a pre-computed tissue depth distribution, and the facial geometry adaptation to the landmarks. These designs enable the effective generation of multiple realistic faces with varying facial features, skin textures, and adiposity levels. Our results demonstrate that the proposed pipeline can effectively generate faces that closely conform to the given skull. This work provides a new tool for law enforcement agencies to generate and explore possible faces from skeletal remains, which greatly facilitates forensic identification when other identifying information is unavailable.
Limitations.
A limitation of our current approach is the constrained complexity of the tissue depth model due to the limited size of the available skull-face dataset. The high cost and prohibitive nature of collecting a large dataset have restricted our ability to train more sophisticated models. We will explore potential synthetic data generation techniques to enhance the dataset size and diversity, thereby enabling the exploration of more complex models for facial reconstruction.
Appendix 0.A Implementation Details
0.A.1 Text Prompt
We aim to synthesize 2D portraits that are depicted from the frontal view without occlusion. Hence, we designed the following prompt template to synthesize 2D portraits as initial faces,
@@ 0.A.1 Text Prompt
To reconstruct the face of Robert the Bruce, King of Scotts, from the scanned skull. We used the following text prompt to generate the initial faces from his historical background,
prompt: A photo of a king’s face. A Scottish male with a mustache, long hair, crown, and royal attire. Looking at the camera. Realistic portrait.
Table 3: Attributes used for initial face generation to match the biological profile of the skull.
@@
0.A.2 Network Weights and Hyper-parameters
0.A.3 Dataset Processing
0.A.4 Pipeline Running Time
@@
Appendix 0.B Analysis on Tissue Depth Distribution
@@
Appendix 0.C More Analysis on Face Adaptation Module
0.C.1 Qualitative Results of Loss Functions
0.C.2 User Study: Convergence Quality of Final Facial Reconstruction from different initial faces
Fig. 13 shows two examples of user study questions. For each question, we randomly synthesized four faces from four different portraits and used the proposed Anatomy-guided Face Adaptation module to adapt to the query skull. Faces A, B, and C are final reconstructions of one skull. Face D is the final reconstruction of the other skull. Faces are shuffled in the user study. Users are asked to compare the reconstructed faces and select the unique one reconstructed from the other skull.
We collected 500 valid responses from 50 valid users. For each question, we count 1 score so the overall score is 10 for 10 questions. Fig. 14 shows the user study results, where the average correctness is $91.2\%$. The result of the user study demonstrates that (1) our model can adapt different initial faces and converge to face shapes that are closely related to the query skull, resulting in distinct reconstructions for different skulls. (2) The reconstructed faces from one skull are not identical, they still preserved subtle geometric details inherited from the initial faces.
@@ 0.C.2 User Study: Convergence Quality of Final Facial Reconstruction from different initial faces
Appendix 0.D Detailed Results of Regional Control
@@
Figure 17: Qualitative comparisons with existing methods. (a) is the input skull. (b-e) show the results of HF-GGR [22], Pix2Pix [20], CycleGAN [58], and CGAN-CFG [56]. (f) show our reconstructed face with texture and geometry. (g) is the CT scanned groundtruth.
Appendix 0.E Extensive Comparisons with Existing Methods

### v2

@@ IV-A Dataset and Error Metric
$E=\frac{1}{d|G|}\sum_{v\in G}~{}\underset{u\in F}{\min}~{}\|v-u\|_{2},$ (7)
@@
IV-B Initial Face Generation
@@
IV-C Evaluation of Tissue Depth Distribution
IV-C1 TDD-Global Model
@@
In the second experiment, we modified the face shapes of celebrities to show the ability of 3D face adjustment by the tissue depth model. Fig. 6 shows the results of three celebrities: Christian Bale, Adele Adkins, and Ling Jia. Using a 2D portrait, the DECA model generates a 3D face of a celebrity, $F_{0}$, with known facial landmarks $\mathcal{Q}^{f}$, thanks to DECA’s consistent face topology. To illustrate the transition of thin to round faces, we sample tissue depths with $C^{\prime}<0$ in the TDD-Global model ($TG$) and gradually increase $C^{\prime}$ to sample different tissue depths $\mathcal{D}={d_{i}}$. The process is reversed to demonstrate the opposite transition. We compute the skull landmarks as $\mathcal{P}^{s}=\mathcal{Q}^{f}-\mathcal{D}$. Next, we vary $C^{\prime}$ to sample different valid tissue depth combinations (e.g., from thin to round faces), where $\mathcal{D}^{\prime}=TG(C^{\prime})$, and then calculate the target facial landmarks $\mathcal{P}^{f}=\mathcal{P}^{s}+\mathcal{D}^{\prime}$ based on these skull landmarks. Finally, the proposed anatomy-guided face adaptation aligns the 3D face $F_{0}$ to the new facial landmarks $\mathcal{P}^{f}$, producing the modified face $F^{\prime}$. We then compare $F^{\prime}$ with another image of the same person in a different face shape.
@@ IV-C2 TDD-Regional Model
Figure 8: Validating the influence of initial faces on the final results. (a) are the query skulls and facial landmarks. (b) is the synthesized portrait image. (c) are the initial 3D faces with texture and the reconstruction errors. (d) are our facial reconstructions, which and fit the landmark constraints well. The final reconstruction errors are similar: 0.899% (top), 0.925% (middle), and 1.19% (bottom). They are close to the GT face (e). The bottom row shows the facial reconstruction from the FLAME model’s average face. Although our face adaptation module can edit the reconstructed geometry to the GT face, without the proposed initial face generation, using average face lacks geometric details and texture.
@@
Figure 10: Extensive qualitative comparisons with existing methods. (a) is the input skull. (b-e) show the results of HF-GGR [28], Pix2Pix [34], CycleGAN [35], and CGAN-CFG [10]. (f) show our reconstructed face with texture and geometry. (g) is the CT scanned groundtruth.
@@
IV-D Evaluation of Skull-to-Face Reconstruction
IV-D1 Influence of Initial Faces on Final Results
We chose the average tissue depths to compute the extended landmarks as facial constraints. Fig. 8 shows that with face refinement, our face adaptation model can effectively modify the face shape to align with the facial landmarks. With facial landmark constraints, the two initial faces, after adaptation, converge to a similar face geometry. Their geometries still exhibit subtle differences inherited from the initial faces (see error maps in (d)). We also use the FLAME model’s average face as the initial face for adaptation. The bottom row of Fig. 8 presents the facial reconstruction results, with final reconstruction errors of 0.899% (top, face candidate 1), 0.925% (middle, face candidate 2), and 1.19% (bottom, average face). Although the reconstructed geometry from the average face is slightly less accurate than the two candidates, it remains close to the ground truth. This demonstrates that our face adaptation module effectively edits the face to fit geometric constraints. Notably, without the proposed initial face generation, the reconstruction from the average face lacks geometric detail and texture. Our initial face generation serves as a powerful visualization tool, providing users with a textured face from text prompts (several keywords or phrases), which offers valuable reference information.
@@ IV-D1 Influence of Initial Faces on Final Results
We also conducted a user study to assess the convergence quality of the final facial reconstruction results from different initial faces. For each question, four initial faces are randomly synthesized. Three of them are used to reconstruct faces from Skull A, and the other one is used to reconstruct a face from Skull B. Users are required to identify the face geometries amongst the reconstructed faces and select the unique face that is reconstructed from Skull B.
For each question, we randomly synthesized four faces from four different portraits and used the proposed Anatomy-guided Face Adaptation module to adapt to the query skull. Faces A, B, and C are final reconstructions of one skull. Face D is the final reconstruction of the other skull. Faces are shuffled in the user study. Users are asked to compare the reconstructed faces and select the unique one reconstructed from the other skull. Fig. 15 in supplementary shows two examples of user study questions.
We collected 500 valid responses from 50 valid users. For each question, we count 1 score so the overall score is 10 for 10 questions. The average correctness is $91.2\%$. The result of the user study demonstrates that (1) our model can adapt different initial faces and converge to face shapes that are closely related to the query skull, resulting in distinct reconstructions for different skulls. (2) The reconstructed faces from one skull are not identical, they still preserved subtle geometric details inherited from the initial faces. Detailed user study result is shown in the supplementary Fig. 14.
The high average correctness of this user study indicates that our model can adapt different initial faces and converge to a face shape space that is related to the query skull, which is distinct from the reconstructed faces that are from the other skulls.
TABLE I: Quantitative comparisons of facial reconstructions with existing methods. The distance between two ears of the groundtruth skin is normalized to 1.
@@
IV-D2 Comparisons with SOTA Facial Reconstruction Methods
@@
Tab. I shows the quantitative comparisons of facial reconstructions. We selected three values of $C$—representing thinner, average, and rounder face types—for all test cases and reported both the average and best reconstruction accuracies. Our method used synthesized portraits as initial faces. Ours (avg. tissue) used the average tissue depths from TDD-Global to compute the facial landmarks and then adapted the initial faces to the final results. Ours (best tissue) utilized three types of tissue depths thin, normal, and fat from TDD-Global as shown in Fig. 3. TDD-Regional is provided to allow users to fine-tune the reconstructed face, while TDD-Global was used here for comparison, as it samples a set of reasonable tissue depth values. We selected the smallest reconstruction result as the final face. Our method outperforms existing methods in all metrics, largely due to the effective adaptation of our loss function to geometric constraints. The CT scan resolution is 0.25mm/pixel. The average of the reconstruction error is 3.464 millimeters.
@@
To evaluate the distribution bias of training/test set, we used 5-fold cross-validation to split the dataset into training (80%) and testing (20%). The average reconstruction errors on the five testing sets are: 1.761%, 1.454%, 1.828%, 1.575%, and 1.665%. The results are consistent with Table I.
IV-D3 Validation on Real-world Facial Reconstruction
@@
V Conclusions
We proposed a novel anatomy-guided 3D face reconstruction approach that can reconstruct faces from a given skull using both biological profiles of the skull and the statistical tissue depth information defined on physiological landmarks. Our pipeline consists of three main components: biological profile-based 3D face initialization, facial landmark prediction using pre-computed tissue depth distributions, and facial geometry adaptation to these landmarks. These components enable the effective generation of multiple realistic facial variations, encompassing diverse features, skin textures, and adiposity levels. Our results demonstrate that the proposed pipeline can reliably generate faces that closely conform to the given skull. This work introduces a valuable tool for applications such as forensic investigations and archaeological reconstructions, allowing users to generate and explore potential faces from skeletal remains.
Limitations. A limitation of our current approach is the constrained complexity of the tissue depth model due to the limited size of the available skull-face dataset. The high cost and prohibitive nature of collecting a large dataset have restricted our ability to train more sophisticated models. We will explore potential synthetic data generation techniques to enhance the dataset size and diversity, thereby enabling the exploration of more complex models for facial reconstruction.
Acknowledgments
We extend our gratitude to Huijun Han and Zhuowen Shen for their assistance in running experiments, and to Qihua Dong and Xianhe Jiao for their efforts in data collection and cleaning.
Appendix A Demo Videos
In addition to this supplementary material, we refer the reader to the supplementary videos for a more comprehensive explanation and additional results.
  1. 1.
    demo1-pipeline.mp4 shows the end-to-end pipeline of the facial reconstruction from a skull. This video shows animation to support the first and the second contributions.
  2. 2.
    demo2-landmark.mp4 shows the demo of the third contribution, the animations of 3D facial editing, globally and in local regions.
  3. 3.
    demo3-comparisons.mp4 shows the comparison between our method and existing methods.
  4. 4.
    demo4-king-face.mp4 shows a real-world facial reconstruction example, which uses the skull of Robert the Bruce, King of Scots from 1306 to 1329.
Appendix B Implementation Details
B-A Text Prompt
We aim to synthesize 2D portraits that are depicted from the frontal view without occlusion. Hence, for the automatic face reconstruction pipeline, we designed the following prompt template to synthesize 2D portraits as initial faces,
@@ B-A Text Prompt
To reconstruct the face of Robert the Bruce, King of Scots, from the scanned skull, we searched for his related information [65], and composed the following text prompt to depict his appearance and infer his initial face:
prompt: A photo of a King of Scots in 14 century. A powerful and strong Scottish male. He has a large head dark brown hair and brown eyes. He also has a muscular neck and strong frame, preparing him for the brutality of medieval warfare. A crown and royal attire. Looking at the camera. Realistic portrait.
B-B Network Weights and Hyper-parameters
B-C Dataset Processing
B-D Pipeline Running Time
TABLE II: Attributes used for initial face generation to match the biological profile of the skull.
@@ B-D Pipeline Running Time
Appendix C Ablation Study on Face Adaptation
We conducted an ablation study to evaluate the impact of various loss functions via three settings: (1) using only $\mathcal{L}_{lmk}$, (2) using $\mathcal{L}_{lmk}$ + $\mathcal{L}_{proj}$, and (3) our complete loss. Tab. III reports the average reconstruction errors. We computed the errors in the physical metric (millimeter) by the average face width (20cm). The projection loss $\mathcal{L}_{proj}$ aids in reducing the distance between the reconstructed face and landmark constraints. The symmetric loss $\mathcal{L}_{sym}$ enhances the face realism by promoting symmetry. The second and third rows show the face adaptation results of the same skull, using two different initial faces.
TABLE III: Reconstruction errors before vs after adaptation and loss functions.
Ablation Study Mean Recon. Mean Recon.
(%) (millimeter)
1. Before adaptation (from DECA) 2.631 5.262
2. w/ $\mathcal{L}_{proj}$ 2.273 4.547
3. w/ $\mathcal{L}_{lmk}$ 1.836 3.672
4. w/ $\mathcal{L}_{lmk}+\mathcal{L}_{proj}$ 1.758 3.576
5. After adaptation 1.732 3.464
($\mathcal{L}_{lmk}+\mathcal{L}_{proj}+\mathcal{L}_{sym}$)
@@ Appendix C Ablation Study on Face Adaptation
Appendix D User Study
Fig. 14 shows the detailed user study result. Fig. 15 shows two examples of user study questions.
Appendix E Analysis on Tissue Depth Distribution
@@
Appendix F Detailed Results of Regional Control